# Modeling the amplification of epidemic spread by misinformed populations

### v3

@@ Methods
SMIR model.
For both the ordinary and misinformed subpopulations, the Susceptible Misinformed Infected Recovered (SMIR) model replicates the standard SIR compartments, denoted as $S_{O}$/$I_{O}$/$R_{O}$ and $S_{M}$/$I_{M}$/$R_{M}$, respectively. While SMIR adopts distinct transmission parameters for the misinformed and ordinary groups ($\beta_{M}$ and $\beta_{O}$, respectively), the mean-field approach assumes that infected individuals from either group ($I_{O}+I_{M}$) can potentially infect anyone in the susceptible populations. The mean-field model is governed by the following system of equations:
To account for homophily, we modify the term $I_{O}+I_{M}$ in Eq. 1 to account for increased (decreased) contacts within (across) groups, according to the parameter $\alpha\geq 0.5$. When homophily does not play a role ($\alpha=0.5$), there is an equal probability of interacting with either subpopulation’s infected group. We thus obtain:
[…]
We initialize the simulation by randomly marking $\frac{\epsilon}{2}$ nodes as infected in both the misinformed and ordinary groups, where $\epsilon=0.001$.
@@
Twitter posts in the CoVaxxy dataset22 were collected in real time via the stream/filter endpoint of the Twitter Application Programming Interface (API). To capture the online discourse surrounding COVID-19 vaccines in English, a comprehensive set of English-language keywords was carefully curated. Beginning with the initial seeds of “covid” and “vaccine,” a snowball sampling technique23 was used to identify co-occurring relevant keywords in December 202022. The resulting list contained almost 80 keywords, available online21. To confirm the relevance of the collected tweets to the topic of vaccines, we examined the coverage obtained by incrementally adding keywords, starting with the most common ones. Over 90% of the tweets in 2021 contained at least one of the three most common keywords: “vaccine,” “vaccination,” or “vaccinate.” To infer the location of accounts, we used the Carmen Python library24 that leverages self-reported location metadata within user profiles (embedded in tweets). As an account’s location may change over time (captured across multiple tweets), we utilize the most recent location. We geolocate 2,047,800 users residing in all 50 U.S. states, who shared a total of 25,806,856 tweets by mapping self-reported locations to U.S. counties. The information network is constructed from accounts in 341 counties that contain more than 200 Twitter users each. Political alignment is estimated using a third-party list of annotated news sources44, 43. It is averaged across all the sources shared by each account. Nodes with an estimated alignment greater (smaller) than zero are considered Republican (Democrat). We infer the political alignment of some additional accounts, who did not share links to news sources, using a label-propagation algorithm18 on the retweet network. If all of a node’s neighbors have political alignment scores, its score is estimated using the weighted average of its neighbors, with weights based on retweets. The process is iterated until each node without a score has at least one neighbor without a score. Misinformation is defined at the source level. Tweets containing links to articles from a list of low-credibility sources compiled by NewsGuard (score below 60) are labeled as spreading misinformation. This approach is common practice and has been validated in the literature39, 28, 8, 31, 48.
@@
To construct edges in our contact network, we utilize SafeGraph cell-phone mobility data55, which contains information on the number of people residing in over 200K Census-Block-Groups (CBG) who visited 4.3M Points-of-Interest (POI) in the United States. This data has been widely employed to study human mobility patterns during the COVID pandemic. We used the average daily number of individuals moving during 2019, as a reference for business-as-usual mobility, and aggregated all CBGs and POIs at the county level. This aggregation results in a county-by-county matrix $L$, where each element $L_{xy}$ represents the average daily number of individuals in county $x$ moving to county $y$ or vice versa. We then normalized $L_{xy}$ to obtain the average probability of individuals in counties $x$ and $y$ coming into contact, and multiplied by the total number of edges to obtain the expected number of connections between individuals in counties $x$ and $y$: $E_{xy}=\frac{L_{xy}}{\sum_{x^{\prime},y^{\prime}}L_{x^{\prime}y^{\prime}}}\frac{\bar{k}N}{2}$ where the sum is over all county pairs and $\frac{\bar{k}N}{2}$ is the total number of edges. Next, we create a physical contact network with $N$ nodes by following a procedure akin to a stochastic block model30 used to generate networks with localized communities. For each pair of distinct locations $x$ and $y$, we draw $E_{xy}$ edges between random pairs of nodes in $x$ and $y$. Additionally, we draw $E_{xx}$ edges among random pairs of individuals within the same location $x$, representing homogeneous mixing within each county. At the end of the process, the network has the target average degree $\bar{k}$. We use $\bar{k}=25$ and show how this parameter affects the infections in Supplementary Information.
@@
Agent-based SMIR simulations are initiated by randomly selecting 100 misinformed nodes and designating them as infected. The disease spreading dynamics are then simulated for 100 steps, which correspond to days. To align with COVID-19 dynamics, we utilize the CDC’s recommended quarantine period of 5 days as our recovery period15 ($\gamma=0.2$). Each simulation is repeated ten times, and the average outcome is reported.
@@
Code and data are available in a public repository (github.iu.edu/NaN-team/bounding-misinfo-impact-on-disease-spread).
@@ Acknowledgments
The initial concept for this research was developed by YA, SF, AF, FM. The final study design was developed collectively by all authors. Data collection was led by MRD as part of the CoVaxxy project 21. All mean-field modeling was conducted by MRD. FP developed the early code for the agent-based simulation, which was subsequently refined and expanded by MRD for publication. Analysis was conducted primarily by MRD, with help from FP, and guidance from YA, SF, AF, FM. The SMIR model was developed collaboratively by all authors. Visualizations were created by MRD, FP, and FM with input from all authors. The first draft was written by MRD, with revisions from all authors. FM oversaw the progression of the study.
@@ Appendix 1 Mean-field SMIR model
Figure S1: Reducing the transmission parameter $\beta_{O}$ decreases the severity of the epidemic. We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values, and (c) the total proportion of the population infected as $\beta_{O}$ varies. Here we do not consider the role of misinformation or homophily.
@@
Figure S2: Effects of varying the recovery rate. We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values, and (c) the total proportion of the population infected as a function of the number of days to recover, $\tau$. Here we do not consider the role of misinformation or homophily.
@@
We now explore the effect of the recovery rate, again in the scenario with no misinformation or homophily, by setting $\beta_{O}=0.3$ and varying the recovery period $\tau=1/\gamma$ between 1 and 20 days. Fig. S2 shows that when $\tau<4$, $R_{0}<1$ and the disease does not reach epidemic proportions. At this level, the epidemic takes a long time to reach its peak ($\approx$80 days). Increasing $\tau$ means that individuals remain infected longer, so the population gets infected faster and the peak infection is reached more rapidly. We set $\tau=5$ ($\gamma=0.2$) in the main text to align with quarantine recommendations from the CDC15. With this value and $\beta_{O}=0.3$, the dynamics are at epidemic levels ($R_{0}=1.5$).
@@ Appendix 2 Robustness analyses
We also investigate the impact of the contact network density ($\bar{k}$) on infection dynamics, as illustrated in Fig. S4 for $5\leq\bar{k}\leq 25$. We consider this range because $\bar{k}=25$ represents pre-pandemic daily social contacts while $\bar{k}=5$ represents COVID-19 lockdown conditions32. As expected, Fig. S4(a) demonstrates that higher $\bar{k}$ leads to increased infections through the population, since the higher contact density provides more opportunities for transmission. But while a larger percentage of the overall population is infected, the relative effect of misinformed individuals decreases. This is because, at higher $\bar{k}$ values, the infected population is already substantial even in the low-misinformation ($\phi=20$) baseline. The combined effect of these two opposing trends, as shown in Fig. S4(b), is that the additional percentage of infected individuals relative to the $\phi=20$ scenario reaches a maximum for some intermediate $\bar{k}$. Fig. S4(b) also shows that, consistent with our primary findings, larger $\phi$ (misinformed resilience) results in a smaller infected population. In our main analysis we focus on $\bar{k}=25$ and model an effective reduction of contacts, resulting for example from social distancing or lockdowns, by decreasing the $p$ parameter.

### v4

@@ Methods
Twitter posts in the CoVaxxy dataset23 were collected in real time via the stream/filter endpoint of the Twitter Application Programming Interface (API). To capture the online discourse surrounding COVID-19 vaccines in English, a comprehensive set of English-language keywords was carefully curated. Beginning with the initial seeds of “covid” and “vaccine,” a snowball sampling technique24 was used to identify co-occurring relevant keywords in December 202023. The resulting list contained almost 80 keywords, available online22. To confirm the relevance of the collected tweets to the topic of vaccines, we examined the coverage obtained by incrementally adding keywords, starting with the most common ones. Over 90% of the tweets in 2021 contained at least one of the three most common keywords: “vaccine,” “vaccination,” or “vaccinate.” To infer the location of accounts, we used the Carmen Python library25 that leverages self-reported location metadata within user profiles (embedded in tweets). As an account’s location may change over time (captured across multiple tweets), we utilize the most recent location. We geolocate 2,047,800 users residing in all 50 U.S. states, who shared a total of 25,806,856 tweets by mapping self-reported locations to U.S. counties. The information network is constructed from accounts in 341 counties that contain more than 200 Twitter users each. Political alignment is estimated using a third-party list of annotated news sources47, 46. It is averaged across all the sources shared by each account. Nodes with an estimated alignment greater (smaller) than zero are considered Republican (Democrat). We infer the political alignment of some additional accounts, who did not share links to news sources, using a label-propagation algorithm19 on the retweet network. If all of a node’s neighbors have political alignment scores, its score is estimated using the weighted average of its neighbors, with weights based on retweets. The process is iterated until each node without a score has at least one neighbor without a score. Misinformation is defined at the source level. Tweets containing links to articles from a list of low-credibility sources compiled by NewsGuard (score below 60) are labeled as spreading misinformation. This approach is common practice and has been validated in the literature41, 29, 8, 33, 51.
@@
To construct edges in our contact network, we utilize SafeGraph cell-phone mobility data61, which contains information on the number of people residing in over 200K Census-Block-Groups (CBG) who visited 4.3M Points-of-Interest (POI) in the United States. This data has been widely employed to study human mobility patterns during the COVID pandemic. We used the average daily number of individuals moving during 2019, as a reference for business-as-usual mobility, and aggregated all CBGs and POIs at the county level. This aggregation results in a county-by-county matrix $L$, where each element $L_{xy}$ represents the average daily number of individuals in county $x$ moving to county $y$ or vice versa. We then normalized $L_{xy}$ to obtain the average probability of individuals in counties $x$ and $y$ coming into contact, and multiplied by the total number of edges to obtain the expected number of connections between individuals in counties $x$ and $y$: $E_{xy}=\frac{L_{xy}}{\sum_{x^{\prime},y^{\prime}}L_{x^{\prime}y^{\prime}}}\frac{\bar{k}N}{2}$ where the sum is over all county pairs and $\frac{\bar{k}N}{2}$ is the total number of edges. Next, we create a physical contact network with $N$ nodes by following a procedure akin to a stochastic block model32 used to generate networks with localized communities. For each pair of distinct locations $x$ and $y$, we draw $E_{xy}$ edges between random pairs of nodes in $x$ and $y$. Additionally, we draw $E_{xx}$ edges among random pairs of individuals within the same location $x$, representing homogeneous mixing within each county. At the end of the process, the network has the target average degree $\bar{k}$. We use $\bar{k}=25$ and show how this parameter affects the infections in Supplementary Information.
@@
Agent-based SMIR simulations are initiated by randomly selecting 100 misinformed nodes and designating them as infected. The disease spreading dynamics are then simulated for 100 steps, which correspond to days. To align with COVID-19 dynamics, we utilize the CDC’s recommended quarantine period of 5 days as our recovery period16 ($\gamma=0.2$). Each simulation is repeated ten times, and the average outcome is reported.
@@
Code and data are available in a public repository (github.com/osome-iu/bounding-misinfo-impact-on-disease-spread).
@@ Acknowledgments
The initial concept for this research was developed by YA, SF, AF, FM. The final study design was developed collectively by all authors. Data collection was led by MRD as part of the CoVaxxy project 22. All mean-field modeling was conducted by MRD. FP developed the early code for the agent-based simulation, which was subsequently refined and expanded by MRD for publication. Analysis was conducted primarily by MRD, with help from FP, and guidance from YA, SF, AF, FM. The SMIR model was developed collaboratively by all authors. Visualizations were created by MRD, FP, and FM with input from all authors. The first draft was written by MRD, with revisions from all authors. FM oversaw the progression of the study.
@@ Appendix 1 Mean-field SMIR model
For both the ordinary and misinformed subpopulations, the Susceptible Misinformed Infected Recovered (SMIR) model replicates the standard SIR compartments, denoted as $S_{O}$/$I_{O}$/$R_{O}$ and $S_{M}$/$I_{M}$/$R_{M}$, respectively. SMIR adopts distinct transmission parameters for the misinformed ($\beta_{M}$) and ordinary ($\beta_{O}$) groups. (In the agent-based model, these are proportional to $p_{M}$ and $p_{O}$, respectively.) The mean-field approximation assumes that the population is well mixed, ignoring the empirical network structure, and that infected individuals from either group ($I_{O}+I_{M}$) can potentially infect anyone in the susceptible populations. The mean-field model is governed by the following system of equations:
To model homophily, we modify the term $I_{O}+I_{M}$ in Eq. 1 to account for increased (decreased) contacts within (across) groups, according to the parameter $\alpha\geq 0.5$. When homophily does not play a role ($\alpha=0.5$), there is an equal probability of interacting with either subpopulation’s infected group. We thus obtain:
[…]
@@
Figure S1: Reducing the transmission parameter $\beta_{O}$ decreases the severity of the epidemic. We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values for the ordinary population, and (c) the total proportion of the population infected as $\beta_{O}$ varies. In (a), the curve for $\beta_{O}=0.2$ is difficult to see because the proportion of the population infected remains very low throughout the simulation. Here we do not consider the role of misinformation or homophily.
@@
Figure S2: Effects of varying the recovery rate. We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values for the ordinary population, and (c) the total proportion of the population infected as a function of the number of days to recover, $\tau$. Here we do not consider the role of misinformation or homophily.
@@
We now explore the effect of the recovery rate, again in the scenario with no misinformation or homophily, by setting $\beta_{O}=0.3$ and varying the recovery period $\tau=1/\gamma$ between 1 and 20 days. Fig. S2 shows that when $\tau<4$, $R_{0}<1$ and the disease does not reach epidemic proportions. At this level, the epidemic takes a long time to reach its peak ($\approx$80 days). Increasing $\tau$ means that individuals remain infected longer, so the population gets infected faster and the peak infection is reached more rapidly.
In summary, the effects of varying the transmission and recovery parameters are predictable: lower $\beta$ and higher $\gamma$ “flatten the curve” and reduce the negative outcomes of an infection. Based on these explorations we set $\tau=5$ ($\gamma=0.2$) to align with quarantine recommendations from the CDC16. We further set $\beta_{O}=0.3$ such that the basic reproduction number is $R_{0}\geq\beta_{O}/\gamma=1.5$ to ensure epidemic spread within the ordinary population.
Appendix 2 Mean-field analyses
Figure S3: Increasing $\lambda=\beta_{M}/\beta_{O}$ accelerates and amplifies the infection. We use $\gamma=0.2$, and $\mu=0.5$. (a) Overall proportion of the population infected as a function of $\lambda$, for different values of $\beta_{O}$. (b) Proportion of the population infected on each day, for different values of $\lambda$ using $\beta_{O}=0.3$. (c) Extra proportion of the total population that is infected as a function of $\lambda$ ($\beta_{O}=0.3$).
To explore the effects of risky behaviors by misinformed individuals, let us assume two equally-sized subpopulations ($\mu=1/2$) and introduce the scaling factor $\lambda=\beta_{M}/\beta_{O}\geq 1$. Fig. S3(a) illustrates the increasing negative impact of the misinformed subpopulation on the disease-spreading dynamics as $\lambda$ becomes larger. If the ordinary population has very low $\beta_{O}$, $\lambda$ has to be very high for the misinformed population to have an effect. On the other hand, if $\beta_{O}$ is large enough, increasing $\lambda$ leads to a ceiling effect, as $\beta_{M}$ cannot exceed one. The social cost associated with the more risky behaviors by the misinformed group is passed on to the whole network. For example, when $\lambda=3.0$, peak infection for the entire population is reached 38 days earlier than in the $\lambda=1$ case (day 22 vs. 60; Fig. S3(b)), leading to an additional 29.3% of the population becoming infected (Fig. S3(c)).
Figure S4: Increasing the initial proportion $\mu$ of the population that is misinformed, as well as $\beta_{O}$, increases (a) the size of the network that becomes infected and (b) the average $R_{0}$ across the population. Here, we fix $\lambda=100$ to match the ratio used in the main text.
We further explore how the initial size $\mu$ of the misinformed population affects the total proportion of the network that ultimately gets infected. We consider various values of $\beta_{M}$ and $\beta_{O}$ such as to capture the same $\lambda=\beta_{M}/\beta_{O}=p_{M}/p_{O}=100$ as in the main text. When $\mu$ and $\beta_{O}$ are both low, the misinformed population has no impact on the infection (Fig. S4(a)), as $R_{0}<1$ (Fig. S4(b)). However, increasing either parameter crosses the epidemic threshold ($R_{0}>1$) so that a significant portion of the population gets infected.
Appendix 3 Effect of homophily
Figure S5: Homophily in the contact network worsens the infection among misinformed individuals, especially for lower transmission rates. The combined effects of transmission and homophily parameters, $\beta_{O}$ and $\alpha$, are examined with the mean-field approximation when $\lambda=3$, $\gamma=0.2$, and $\mu=0.5$. We plot the proportions of infected individuals in (a) the ordinary population, (b) the misinformed population, and (c) the overall population. The maximum proportion of the overall population infected for each $\beta_{O}$ is marked with a black dot. When the transmission rate is sufficiently high, homophily benefits the entire population but harms the misinformed group.
Fig. S5 illustrates the effects of homophily ($0.5\leq\alpha\leq 1$) for different levels of ordinary transmission ($0.1\leq\beta_{O}\leq 0.4$). For less infectious disease (low $\beta_{O}$), increasing homophily significantly harms the misinformed group (Fig. S5(b)): the infection remains confined within this group. There is no discernible effect on the ordinary population as long as the two groups interact; when they do not ($\alpha=1$), we observe a sharp reduction in infections (Fig. S5(a)). For $0.160.12<\beta_{O}<$ 0.16, peak infection scenarios coincide with intermediate homophily levels, as indicated by the black dots in Fig. S5(c) 39. Under these conditions, while increased homophily decreases infections in the general populace, it significantly worsens outcomes for the misinformed group (compare Fig. S5B and C). As $\beta_{O}$ increases further, while nearly the entire misinformed population becomes infected regardless of $\alpha$ (Fig. S5(b)), high homophily shields the full population (Fig. S5(c)): ordinary individuals have a lower risk of becoming infected through interactions with misinformed individuals. In summary, homophily offers greater protection to the ordinary group by isolating misinformed communities, which suffer a greater disease burden, exacerbating health disparities 49, 53.
@@ Appendix 4 Robustness analyses
Fig. S7 illustrates the impact of the contact network density ($\bar{k}$) on infection dynamics, for $5\leq\bar{k}\leq 25$. We consider this range because $\bar{k}=25$ represents pre-pandemic daily social contacts while $\bar{k}=5$ represents COVID-19 lockdown conditions34. As expected, Fig. S7(a) demonstrates that higher $\bar{k}$ leads to increased infections through the population, since the higher contact density provides more opportunities for transmission. But while a larger percentage of the overall population is infected, the relative effect of misinformed individuals decreases. This is because, at higher $\bar{k}$ values, the infected population is already substantial even in the low-misinformation ($\phi=20$) baseline. The combined effect of these two opposing trends, as shown in Fig. S7(b), is that the additional percentage of infected individuals relative to the $\phi=20$ scenario reaches a maximum for some intermediate $\bar{k}$. Fig. S7(b) also shows that, consistent with our primary findings, increasing $\phi$ (misinformed resilience) decreases the infected population. In our main analysis we focus on $\bar{k}=25$ and model an effective reduction of contacts by decreasing the $p$ parameter.